A sum of R1 000 is due in five months' time from now, at a simple interest rate of 12% per annum. The present value of the sum is

1.     R952,38.

2.     R567,43.

3.     R1 050,00.

4.     R625,00.

5.     none of the above


1
Expert's answer
2021-05-19T17:56:07-0400

"principal\times Rate\times Time+principal=1000"

"principal\times 0.12\times 5/12+principal=1000"


"principal=\frac{1000}{1.05}=952.38"


Answer is option 1 of R 952.38